### v2

# LLM-as-a-qualitative-judge: automating error analysis in natural language generation

Abstract
Prompting large language models (LLMs) to evaluate generated text, known as LLM-as-a-judge, has become a standard evaluation approach in natural language generation (NLG), but is primarily used as a quantitative tool, i.e. with numerical scores as main outputs. In this work, we propose LLM-as-a-qualitative-judge, an LLM-based evaluation approach with the main output being a structured report of common issue types in the NLG system outputs. Our approach is targeted at providing developers with meaningful insights on what improvements can be done to a given NLG system and consists of two main steps, namely open-ended per-instance issue analysis and clustering of the discovered issues using an intuitive cumulative algorithm. We also introduce a strategy for evaluating the proposed approach, coupled with $\sim$300 annotations of issues in instances from 12 NLG datasets. Our results show that LLM-as-a-qualitative-judge correctly recognizes instance-specific issues in 2/3 cases and is capable of producing error type reports resembling the reports composed by human annotators. Our code and data are publicly available [^1] Code & data: https://github.com/tunde-ajayi/llm-as-a-qualitative-judge.
LLM-as-a-qualitative-judge:
automating error analysis in natural language generation
[…]
1 Introduction
Prompting large language models (LLMs) to output evaluation scores, known as LLM-as-a-judge (Zheng et al., 2023a), has become a standard approach for evaluating performance in natural language generation (NLG) tasks. In contrast to classic statistical measures such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), or METEOR (Banerjee and Lavie, 2005), which primarily rely on surface-level lexical overlap, LLM-as-a-judge evaluates text based on semantic understanding, allowing it to better handle diverse phrasings that convey equivalent meanings. Consequently, it shows stronger alignment with human judgment in various tasks, including machine translation (Kocmi and Federmann, 2023), summarization (Clark et al., 2023), or open-ended instruction following (Ye et al., 2024), especially with strong recent LLMs.
Figure 1: Issue types reports for two datasets composed by the proposed LLM-as-a-qualitative-judge (GPT-4o) and by a human annotator. All steps of analysis performed by GPT-4o, including error types formulation and error grouping. The full generated report also includes comprehensive error type descriptions, omitted here due to the space limit. Cnt represents issue type counts.
Figure 2: Illustration of the proposed approach and evaluation methodology.
Recent works propose various extensions of the LLM-as-a-judge approach, including pairwise model comparison (Zheng et al., 2023b), finetuning LLMs for evaluation (Kim et al., 2024b, c), multi-criteria evaluation (Liang et al., 2023; Fu et al., 2024), the dynamic selection of evaluation criteria (Ye et al., 2024), or even using per-instance evaluation checklists (Cook et al., 2024; Kim et al., 2025). A common technique to improve LLM-as-a-judge is to ask a model to output an explanation for the predicted score(s).
However, even with the extensions listed above, LLM-as-a-judge remains primarily a quantitative evaluation tool, i.e., the final result used by researchers and practitioners is quantitative evaluation scores. At the same time, language generation is a complex multi-faceted task with a vast space of potential issues, including in various aspects of generated texts (grammaticality, factuality, logical coherence, etc), in preprocessing of the input data and postprocessing of the NLG outputs, or even with user requests. An effective and commonly used strategy for spotting such issues is a manual qualitative error analysis of a subset of predictions, which allows developers to identify artifacts, fix system issues, and detect flaws in quantitative evaluation. Yet this analysis is often skipped in practice, due to overreliance of developers on quantitative metrics, as well as high demand in terms of time and effort needed to conduct such analysis.
In this work, we introduce LLM-as-a-qualitative-judge, a novel approach which automates error analysis, with the main output being a a structured report aggregating the common qualitative error types in the NLG outputs for a given dataset. The two key steps of LLM-as-a-qualitative-judge are (1) open-ended per-instance error analysis and (2) clustering of the discovered error types. Per-instance analysis implies prompting an LLM to detect an issue in the given NLG system output, where an issue may be arbitrary, i.e. we do not provide any predefined set of possible issues. For error clustering, we propose an intuitive and effective algorithm which resembles how humans solve the corresponding task. The examples of the generated reports are presented in Figure 1, and a high-level illustration of the proposed approach is presented in Figure 2.
To evaluate the proposed LLM-as-a-qualitative-judge, we manually annotate error types in a diverse set of 12 generative datasets (7 tasks and 6 domains for one of the tasks) and introduce an LLM-as-a-judge-based strategy for comparing the errors detected by an LLM and by a human expert. We also manually annotate potential clustering of errors in each dataset, which allows us to evaluate the automatically produced clustering, by calculating its agreement with the human-annotated clustering.
To summarize, our contributions are three-fold:
  • We introduce an LLM-as-a-qualitative-judge, a novel approach for LLM-based evaluation, outputting a structured report of common error types in a dataset;
  • We collect a set of $\sim$300 manual annotations of open-ended issues in the instances coming from 12 diverse NLG datasets, as well as the manual annotations of their per-dataset clustering;
  • We introduce a strategy for evaluating LLM-as-a-qualitative-judge and show that LLM-as-a-qualitative-judge is capable of producing error type reports which resemble the reports produced by humans.
We hope that the proposed LLM-as-a-qualitative-judge approach will reduce the time and effort required for issue analysis and will help practitioners to more easily spot issues in their NLG pipelines.
2 Proposed approach
The main goal of our proposed approach, LLM-as-a-qualitative-judge, is to provide a developer with a structured report of the main types of issues (and their counts) in the outputs of a given NLG system for a given dataset. In the rest of the work, we use terms issues, errors, or failures interchangeably to denote any problems which may occur in the NLG outputs. Examples of such problems include (but are not limited to) unfinished generation due to reaching the maximum new tokens limit, an error in one of the reasoning steps, a problem with the retrieved documents in retrieval-augmented generation, or an error in evaluation due to the use of an inappropriate metric. We do not employ any predefined set of possible issues, and use the term open-ended issue analysis to refer to the problem of detecting arbitrary issues in NLG outputs.
For the purposes of our algorithm, the dataset consists of instances, each represented by a user input (a string containing a task instruction and the input data), a ground truth response (a string defining a correct answer), and a generated response (a final output of the NLG system). Each instance can be optionally augmented with other fields, e.g., the intermediate outputs of an NLG system such as retrieved documents in retrieval-augmented generation, or additional information on the NLG system, e.g., a definition of a task metric. The LLM-as-a-qualitative-judge algorithm is summarized in Algorithm 1, illustrated in Figure 2 and described step-by-step below.
Preliminary step: detecting examples with errors.
Our algorithm focuses only on the instances from the dataset with any sort of issues. We rely on the task-specific quantitative metric to select such instances, i.e., instances which did not get high scores in the quantitative evaluation.
Step 1: per-instance analysis.
For each instance, we prompt an LLM to identify “one, most important, specific, clearly visible issue”, provided with a user input, a ground truth response, a generated response, and optionally other fields as described above. We prompt an LLM to output a detailed analysis of a given instance, followed by a special separator and a final 1–2 sentence description of an identified issue, which is referred to as a per-instance issue explanation in the following steps of the approach. The particular prompt used for per-instance analysis is presented in App. Figure 5.
Algorithm 1 LLM-as-a-qualitative-judge
1:a list of user inputs $U$, a list of ground truth responses $R^{\mathrm{gt}}$, a list of generated responses $R$ — all of length $N$;
2:a report $C$ listing issue types and their counts;
a list $A$ of per-instance issue explanations
[…]
5:for $i=1,\dots,N$ do
6:  // per-instance analysis
7:  $A[i]\leftarrow\mathrm{LLM}(\mathrm{prompt}_{\mathrm{analysis}};U[i],R^{\mathrm{gt}}[i],R[i])$
8:  // $A[i]$ is a string containing issue explanation
9:  // report generation
10:  if $i>1$ then
11:   // an existing issue type or a new one?
12:   $K\leftarrow\mathrm{LLM}(\mathrm{prompt}_{\mathrm{decision}};A[i],C)$
13:   // $K\in\{1,\dots,|C|,\mathrm{None}\}$
14:  else
15:   $K\leftarrow\mathrm{None}$ // first step is always new issue type
16:  if $K$ is None then
17:   // create a new issue type
18:   $E\leftarrow\mathrm{LLM}(\mathrm{prompt}_{\mathrm{new\_type}};A[i])$
19:   // $E$ is a dictionary containing a short issue name and an issue description
[…]
21:   $C.\mathrm{append}(E)$
22:  else
23:   // augment an existing issue type
[…]
25:return $C$, $A$
Step 2: issue clustering.
The second step in LLM-as-a-qualitative-judge consists of clustering issues discovered in the first step and forming a final report of main issue types based on the clustering results. This can be, in principle, done with any clustering algorithm, e.g., k-means with BERT-based embeddings (Devlin et al., 2019) or directly prompting a strong LLM to output clustering (Viswanathan et al., 2024), provided with clustering inputs in a single prompt. In the experiments, we demonstrate the downsides of these approaches, e.g., classic approaches perform poorly on our data, and clustering with direct prompting fails for larger datasets, weaker LLMs, and does not ensure the structural correctness of the generated report.
Inspired by how humans would cluster issues, we propose an intuitive cumulative issue clustering algorithm. Our clustering algorithm goes through instances one-by-one and gradually builds the issue types report. For each instance, we provide an LLM with the current report and the per-instance issue explanation, and prompt the LLM to decide whether this issue explanation can be attributed to one of the already discovered issue types (clusters) or it should form a new issue type (cluster). In the former case, we augment the counter of the corresponding issue type by one. In the latter case, we also prompt an LLM to formulate a short name and a 1–2 sentence description of a new error type, based on the per-instance issue description. In particular, we instruct an LLM to formulate “a fine-grained issue type that can be generalized to other instances”. The new issue type is then added to the report, represented by the generated issue type name, description, and the counter set to one. The first instance in the dataset is always a new issue type. Appendix Figures 7 and 8 show the prompts used for the two described clustering steps.
The final issue types report is composed of issue type names and descriptions, paired with the counts of how many instances were attributed to the corresponding issue type.
3 Evaluation protocol
In this section, we introduce an evaluation protocol to assess the effectiveness of both steps in the LLM-as-a-qualitative-judge approach.
Per-instance analysis.
The core of our evaluation strategy is to collect manual annotations of failure cases for a set of instances from various tasks and domains. For each instance, consisting of a user input, a ground truth response, a generated response, a task metric, and optionally retrieved documents, an annotator’s task is to formulate what the particular issue is in this instance. We then prompt an evaluator LLM to judge whether the issue explanation determined by the LLM-as-a-qualitative-judge for a particular instance matches the issue determined by the human annotator. The outputs from the evaluator LLM are binary and can be accumulated into a per-instance analysis accuracy score.
Task Dataset reference # ex.
Natural Language Generation
Instruction following FLASK Ye et al. (2024) 34
Translation en-ru WMT’22 Kocmi et al. (2022) 38
Long context QA Elitr-Bench Thonet et al. (2025) 26
Semantic parsing PIZZA Arkoudas et al. (2022) 34
Grade school math GSM8K Cobbe et al. (2021) 17
Detoxification ParaDetox Dementieva et al. (2024) 36
Retrieval-augmented QA
Factoid QA in Russian MKQA (ru) Longpre et al. (2021) 29
Biomedical QA BioASQ Krithara et al. (2023) 27
Lifestyle forum QA RobustQA Han et al. (2024); Santhanam et al. (2022) 21
Writing forum QA 13
Search engine queries SearchQA Dunn et al. (2017) 13
Educational QA SyllabusQA (Fernandez et al., 2024) 9
Total 297
Table 1: The statistics of the annotated evaluation data.
The annotation instruction asks to ignore instances which have multiple issues (to avoid ambiguity in per-instance analysis), instances where ground truth labels appear to be wrong, and instances where the annotator’s expertise is insufficient to judge the correctness of the generated answer. We also limit the number of instances with the same issue to not exceed 8 examples per dataset, to ensure the diversity of the final dataset.
Issue clustering.
Human annotation also includes a step of manually clustering issues discovered in per-instance analysis, i.e., specifying cluster indices and cluster names (generalized issue types) for labeled instances. This annotation is then used to compute clustering agreement between the clustering produced by LLM-as-a-qualitative-judge and by a human annotator.
We also consider synthetic data for evaluating clustering: we define a set of possible issue types $e$ and their frequencies $n_{e}$, then prompt GPT-4o to reformulate each issue $e$ in various ways $n_{e}$ times, and then use this data as per-instance analysis for clustering. This allows us to evaluate clustering on larger datasets, i.e., 100-1000 instances.
We evaluate cluster agreement using a Rand index adjusted for chance, or Adjusted Rand Index (ARI) [^2] We use the scikit learn implementation: https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_score.html.. We also evaluate the agreement in error type descriptions, by finding the best possible mapping between clusters found by a human annotator and by LLM-as-a-qualitative-judge, and then prompting an evaluator LLM to judge the semantic equivalence of the corresponding issue type descriptions. This metric is denoted as Semantic Label Consistency (SLC).
Dataset composition.
Table 1 provides the data summary. We consider a diverse pool of 7 generative tasks, and for one of the tasks, namely retrieval-augmented question answering (RA-QA), we consider a diverse pool of 6 domains. For each considered dataset, we manually label failures in up to 40 generations from one of the open-source LLMs (Qwen-2.5-1b, Llama-3.2-1b, Command-R-35b, Vicuna-1.5-13B). All the labeling was performed by the authors of the paper in Google Spreadsheets [^3] https://docs.google.com/spreadsheets. The final dataset comprises 297 instances.
4 Experiments
4.1 Experimental setup
We test per-instance analysis with a range of commercial and open-source LLMs. For issue clustering, we compare the proposed cumulative clustering approach to the direct LLM prompting and classic clustering approaches. All clustering runs operate on the per-instance analysis output by GPT-4o, and clustering results are averaged over 3 runs from different random seeds. We test issue clustering with three LLMs: GPT-4o, Gemini-2-Flash, and Qwen-2.5-7B. Tables 5 and 6 in Appendix list references and licenses of the used models and datasets, respectively.
Figure 3: Examples of per-instance analysis.
Open-source LLMs are run on a single V100 GPU with greedy decoding ($\sim$20 GPU-hours in total). Commercial LLMs are run via API with requesting json output format. Prompts used for all the stages are presented in the Appendix A. Exact task formulations in prompts were adjusted by only using three RA-QA datasets (MKQA (ru), RobustQA Lifestyle and Writing).
For each instance, LLM-as-a-qualitative-judge is provided with a user input, a ground truth response, a generated response, 5 retrieved documents (only for RA-QA), and a short task comment. The task comment describes the task metric (in one sentence), provides a comment on the nature of ground truth responses (either that it is the expected answer or that it is only one of the possible correct answers), and also contains a comment that retrieval-augmented generation (RAG) or Chain-of-Thought (COT) prompting was used, when applicable (6 datasets with RAG and 2 datasets with COT). The used task metrics are binary LLM-as-a-judge (the generated response is accepted or not) or binary Match (outputs True is one of the ground truth answers is contained a substring in the generated response, and False otherwise). Task comments for all datasets are also presented in Appendix E.
For classic clustering approaches, we use the scikit-learn implementation with BERT embeddings and tune hyperparameters as described in App. B. For methods requiring the number of clusters, we set it the same as in the annotator’s data.
Meta-evaluation of an evaluator LLM.
To ensure the reliability of the evaluator LLM, we collected a small meta-evaluation dataset of 50 instances from 4 datasets (MKQA (ru), RobustQA Writing, FLASK, MultiDetox), where the equivalence of the LLM-as-a-qualitative-judge’s and human annotator’s per-instance analysis was judged by a human annotator and can be compared to the evaluator LLM’s verdict. Strong commercial LLMs, such as GPT-4o, Gemini-2.0-Flash, and Claude-3.7-Sonnet, achieved a meta-evaluation accuracy of 85-90% on this dataset, and an open-source Solar-10.7B (Kim et al., 2024a) achieved a meta-accuracy of 60%. In all the experiments, we use claude-3-7-sonnet-20250219 as the evaluator LLM, to avoid using the same LLM for analysis and for evaluation.
Model Accuracy (%)
GPT-4o 66.3
Gemini-2.0-Flash 65.0
Qwen-2.5-32B 68.7
Qwen-2.5-7B 55.5
Qwen-2.5-1.5B 30.7
DeepSeek-R1-Distill-Llama-8B 56.1
Aya-Expanse-8B 42.1
Llama-3.1-8B-Instruct 55.4
Ministral-8B-Instruct-2410 58.1
Table 2: Performance of various LLMs in per-instance analysis. Evaluator LLM: Claude-3-7-Sonnet-20250219.
4.2 Per-instance analysis
Table 2 reports the performance of various LLMs in per-instance analysis. Strongest LLMs, including commercial LLMs and a larger open-source Qwen-2.5-32B achieve an accuracy of 62–67%, i.e. about 2/3 of issues in our dataset were successfully correctly explained by strong models. The accuracy of open-source LLMs is substantially influenced by their size: Qwen-2.5 accuracy raises from 32% to 67% when increasing size from 1.5B to 32B. Various LLMs of 7–8B size demonstrate analysis accuracy of 42–60%.
Figure 4: Examples of confusion matrices visualizing clustering agreement between LLM-as-qualitative-judge-generated and the annotator’s issue types reports. We find the optimal mapping between clusters found by a human annotator and by LLM-as-a-qualitative-judge, and then define a confusion matrix where each cell $(i,j)$ denotes a number of dataset instances allocated into $i$-th annotator’s cluster and $j$-th LLM-as-a-qualitative-judge’s cluster.
Figure 3 shows examples of issue explanations generated by GPT-4o and Qwen-2.5-7B. In the manual inspection of the generated issue explanations, we observed correct explanations for various kinds of issues, and rows 1–3 demonstrate such examples. We also highlighted three common groups of mistakes. The first repeated problem in per-instance analysis is logical issues. In the example in row 4, the issue is that the ground truth response is not contained as a verbatim substring in the model-generated response, which is a definition of a task metric. However, both GPT-4o and Qwen-2.5-7B claim that the failure in the substring match is caused by the model response containing extra generated information. Such an explanation logically contradicts the task metric, i.e., extra content can only increase chances of finding a given substring in the response, but cannot be a reason for its absence.
The second common problem in per-instance analysis is the oversimplification of an issue, especially for more unexpected issues, such as an ambiguous user input or an error in evaluation. In the example in row 5, the issue is an ambiguous user question, i.e., both the ground truth and the generated response are correct and provide two different interpretations of the user question. However, GPT-4o and Qwen-2.5-7B report the over-simplified issue of the generated response not providing the same answer as the ground truth response.
Finally, the third common reason for a per-instance issue explanation not being accepted by the evaluator LLM, is the subjectivity of some issues in the dataset. For example, a human-annotated issue in row 6, “The generation was stopped too early because of the reached maximum new tokens limit”, is evaluated to be not equivalent to the LLM-as-a-qualitative-judge-generated issue “The generated response provides an incomplete overview <…>”. While these two issue explanations are indeed different, they are both correct, and the LLM-generated explanation follows from the annotator’s explanation.
Future work could attempt to alleviate the first two mentioned problems by using stronger reasoning approaches in per-instance analysis, and the third problem by improving an evaluator LLM.
Approach Cluster assignment Cluster descriptions
[…]
GPT-4o
Cumulative 0.14±.05 0.73±.05 0.33±.10 0.70
Direct 0.15±.05 0.63±.04 0.42±.12 0.62
Gemini
Cumulative 0.13±.05 0.70±.02 0.32±.12 0.71
Direct 0.17±.04 0.83±.01 0.42±.11 0.68
Qwen-2.5-7B
Cumulative 0.11±.04 0.50±.07 0.41±.16 0.44
Direct 0.07±.05 0.01±.02 0.32±.11 0.12
K-means 0.05±.05 0.44±.08 n/a n/a
Agglomerative 0.05±.00 0.49±.00 n/a n/a
GMM 0.04±.03 0.41±.05 n/a n/a
HDBSCAN 0.01±.02 0.13±.03 n/a n/a
Table 3: Performance of various approaches and LLMs in issue clustering. Results averaged over 3 runs from different random seeds. Agreement in cluster assignment measured using Adjusted Rank Index (ARI) and in cluster descriptions using LLM-judged Semantic Label Consistency (SLC, evaluated by Claude-3.7-Sonnet). Subscripts ${}_{\text{real}}$ and ${}_{\text{syn}}$ indicate tests on the real and synthetic data respectively. “N/a” indicates the metric is not applicable since classic approaches do not generate cluster names.
4.3 Issue clustering
Table 3 reports performance in issue clustering, for three LLMs. As described in Section 3, we evaluate clustering on both real and synthetic data. We find that clustering via direct prompting performs well for small datasets and strong LLMs, but fails for weaker LLM, e.g. Qwen-2.5-7B, and/or larger datasets. For example, ARI reached by GPT-4o on the synthetic data drops from 1 to 0.05 when increasing the dataset size from 100 to 1000 instances. The proposed cumulative clustering demonstrates greater robustness and reaches high ranges of ARI in all cases. In addition, the proposed cumulative algorithm outputs correctly structured summaries by design, while the structural correctness of clustering with direct prompting is not guaranteed.
Comparing LLMs, we find that Gemini reaches highest scores in both cluster assignment and cluster descriptions generation, followed by GPT-4o and Qwen-2.5-7B. Classic clustering approaches reach rather low values of ARI.
Figure 4 demonstrates examples of confusion matrices for several datasets. Pronounced diagonals and matching cluster names illustrate the strong capabilities of LLM-as-a-qualitative-judge to output comprehensive issue types reports. Due to the inherent subjectivity of clustering task, we observe occasional merging or splitting of annotator’s clusters, e.g. clusters “Wrong topping” and “Wrong variables” were merged by LLM-as-a-qualitative-judge into one cluster “Entity Mislabeling” for the Pizza ordering dataset.
Table 4 demonstrates the non-uniformity of LLM-as-a-qualitative-judge’s performance across datasets, i.e. issue type reports generated by LLM-as-a-qualitative-judge reach high agreement with annotator’s reports for certain datasets and rather low agreement for other datasets. We attribute this effect to the imperfections of per-instance analysis discussed above (which we believe could be further alleviated in future work) and the inherent subjectivity of the issue explanation and clustering tasks. Appendix E present the examples of generated error type reports, per-instance issue explanations, and confusion matrices for all considered datasets. We also provide an archive with experimental results in the project repository [^4] A zip archive with experimental results is available at https://github.com/tunde-ajayi/llm-as-a-qualitative-judge/tree/main/data.
Dataset Per-inst. an. acc. (%) Issue clust. ARI Issue clust. SLC
Semantic parsing 94.1 0.41 0.29
Grade school math 88.2 0.04 0.22
Detoxification 77.8 0.36 0.28
Long-context QA 69.2 0.07 0.50
Translation en-ru 65.8 0.10 0.63
Instruction following 55.9 0.09 0.19
RA-QA: SyllabusQA 77.8 0.16 0.55
RA-QA: MKQA (ru) 75.8 0.17 0.44
RA-QA: BioASQ 66.7 0.08 0.31
RA-QA: SearchQA 38.4 0.15 0.16
RA-QA: Writing 30.7 0.00 0.11
RA-QA: Lifestyle 19.0 0.00 0.32
Table 4: Per-dataset results for GPT-4o as LLM-as-a-qualitative-judge.
5 Related work
Quantitative LLM-based evaluation.
While using commercial LLMs for evaluation remains common practice, one line of work (Kim et al., 2024b, c; Pombal et al., 2025) focuses on tuning open-source LLMs on the synthetically generated evaluation data, to ensure reproducibility of evaluation. Other works improve quantitative LLM-as-a-judge by conducting more fine-grained evaluation, e.g. using multiple evaluation criteria (Liang et al., 2023; Fu et al., 2024) or selecting evaluation criteria individually per instance (Ye et al., 2024; Cook et al., 2024; Kim et al., 2025). Composite evaluation approaches such as FactScore (Min et al., 2023) or RAGChecker (Ru et al., 2024) use LLMs in the intermediate evaluation steps.
Qualitative LLM-based evaluation.
LLM-generated qualitative error explanations are often used to improve the precision of quantitative evaluation (Zeng et al., 2024; Ye et al., 2024) or to explain the assigned quantitative scores to a developer (Xu et al., 2023; Jiang et al., 2024). Such approaches only output per-instance explanations, and a substantial human effort is still needed to read all of them. Certain works (Jiang et al., 2024; Kasner et al., 2024; Perrella et al., 2022) focus on outputting aggregated reports of frequent errors, but with a (limited) predefined error set, i.e. they solve the task of error classification. In contrast to these efforts, LLM-as-a-qualitative-judge outputs an aggregated report of issue types discovered in an open-ended manner, i.e. without any predefined issue set.
Meta-evaluation.
A line of community efforts (Zeng et al., 2024; Lambert et al., 2025; Hada et al., 2024; Bavaresco et al., 2024) is devoted to an important task of meta-evaluating LLM-as-a-judge, i.e. collecting human annotations for various tasks, domains, or languages, and evaluating how closely LLMs mirror human judgments. Certain task-specific datasets (Freitag et al., 2021) can be used to meta-evaluate fine-grained issue detection. Our work further contributes to this direction by the release of a meta-evaluation dataset, containing qualitative issue explanations for 12 datasets from 7 tasks and their per-dataset clustering.
Clustering with LLMs.
Earlier works (Petukhova et al., 2024; Miller and Alexander, 2024) demonstrate advantages of leveraging LLM-derived embeddings in place of traditional TF-IDF or BERT vectors in standard clustering algorithms. More recent works employ LLMs directly to cluster textual data. Viswanathan et al. (2024) instruct a GPT-3.5 model to cluster the provided data given few-shot demonstrations. Huang and He (2024) transform clustering into a two-stage classification task: first prompting an LLM to infer a set of candidate clusters for the dataset, then prompting it to assign the best cluster to each instance. ClusterLLM Zhang et al. (2023) uses an instruction-tuned LLM to guide clustering, i.e., to decide which clusters to merge. In our work, we propose an alternative intuitive approach for LLM-based clustering. Our approach can also be extended in the future with the listed strategies.
6 Conclusion
In this work, we present LLM-as-a-qualitative-judge, a novel approach for generating structured reports summarizing key types of issues in a given NLG system. We hope that this approach will help developers to spot more easily issues and artifacts in their NLG systems. For example, from reports shown in Figure 3, a developer may decide to revisit the training data mixture (left) or a task metric (right). The LLM-as-a-qualitative-judge approach can also be easily extended to compare two NLG systems, by using two separate issue counters (one per dataset) in the cumulative clustering algorithm. Our code and data are available at https://github.com/tunde-ajayi/llm-as-a-qualitative-judge.
Future works could equip LLM-as-a-qualitative-judge with advanced reasoning or agentic pipelines, tune LLMs for issue report generation, and study the approach for a wider set of languages.
Acknowledgments
We greatly appreciate the help of Alexandre Misrahi, Salah Aït-Mokhtar, and Maxime Louis. The project was initiated at the Advanced Language Processing School (ALPS 2025, https://alps.imag.fr).
Appendix A Prompts
You are an expert in analysing the failure cases in natural language generation tasks
You are given a question, ground truth label(s), and the answer generated by an LLM.
The generated answer was not accepted by the automatic evaluation measured with metric {metric info}
Read all these materials and reply what is the particular failure case in this example, i. e. why exactly the generated response was not accepted.
The problem can be in any part of the pipeline, including the question itself or any aspects of the system outputs.
IMPORTANT: identify ONE, MOST IMPORTANT, SPECIFIC, CLEARLY VISIBLE issue in each test case.
Question:
{question}
Ground truth label(s):
{label}
LLM-generated answer:
{answer}
So what is the particular failure in this example? First output a detailed analysis, and then output a final summary of the failure in one or two sentences after a special separator ‘‘Summary:’’.
Figure 5: Prompt used for per-instance analysis. The presented version of the prompt is for text LLM outputs, the prompt can be easily changed if JSON outputs are supported by an LLM. For RA-QA, we also include retrieved documents in the prompt.
You are an expert in analysing the failure cases of natural language generation systems.
You already performed per-example analysis, where for each example, you were given a question, ground truth label(s), and the answer generated by an LLM.
The generated answer in all examples was not accepted by the automatic evaluation.
You already read all these materials and formulated what was the particular failure case in each example, i.e. which part of the pipeline failed so that the generated response was not accepted.
Now your task is to summarize all your per-example analyses into a concise overall summary of failure cases for the given dataset.
Summarize what are the various failure types in this dataset (provide the overall count of each error type and also ids of all examples of each error type).
Please try to be very specific in determining error types, avoid too much generic error types. On the contrary, determine as much as possible FINE GRAINED error types.
Furthermore, provide a comment for each error type explaining the essence of this error type in a bit more details (in the context of this particular dataset).
*** Per-example analysis which you generated before ***
{analysis}
*** per-example analysis ended ***
Summarize all your per-example analyses into a concise overall summary of failure cases.
Generate a json with the only key ‘‘summary’’, and a value is a dict of error types. Each value in this dict (corresponding to one detected error type) is a dictionary with keys ‘‘error_name’’, ‘‘error_description’’, ‘‘indexes’’ (indexes of all examples with this error type), and ‘‘num_examples’’ (overall count of this error type).
Figure 6: Prompt used for summary direct prompting.
You are an expert in analysing the failure cases of natural language generation systems.
You already performed per-example analysis, where for each example, you were given a question, ground truth label(s), and the answer generated by an LLM.
The generated answer in all examples was not accepted by the automatic evaluation.
You already read all these materials and formulated what was the particular failure case in each example, i. e. which part of the pipeline failed so that the generated response was not accepted.
Now your task is to summarize all your per-example analyses into a concise overall summary of failure cases for the given dataset.
You are using a SPECIAL CUMULATIVE ALGORITHM as follows. You are going through examples one by one and accumulate discovered error cases in a special pool. For each example, you check if any of the already discovered error types from the pool fits it, and if so, you assign this error type to this example. If none of already existing error types fit the current example, you create a new error type and add it to a pool.
A pool of already discovered error cases:
{error cases}
Analysis of a current example:
{analysis}
Do you think any of the already discovered error cases fit the current example? If yes, output ‘‘Decision:’’ and a key marking the chosen error case, e. g. ‘‘Decision: type_0’’. Do it only if you are REALLY sure that the chosen error case fits the current example! DO NOT output cluster name. If not, output ‘‘Decision: None’’. DO NOT output anything else.
Figure 7: Prompt used for classifying each instance in the cumulative clustering strategy. The presented version of the prompt is for text LLM outputs, the prompt can be easily changed if JSON outputs are supported by an LLM.
You are an expert in analysing the failure cases of natural language generation systems.
You already performed per-example analysis, where for each example, you were given a question, ground truth label(s), and the answer generated by an LLM.
The generated answer in all examples was not accepted by the automatic evaluation.
You already read all these materials and formulated what was the particular failure case in each example, i. e. which part of the pipeline failed so that the generated response was not accepted.
Now your task is to summarize all your per-example analyses into a concise overall summary of failure cases for the given dataset.
You are using a SPECIAL CUMULATIVE ALGORITHM as follows. You are going through examples one by one and accumulate discovered error cases in a special pool. For each example, you check if any of the already discovered error types from the pool fits it, and if so, you assign this error type to this example. If none of already existing error types fit the current example, you create a new error type and add it to a pool.
A pool of already discovered error cases:
{error cases}
Analysis of a current example:
{analysis}
You decided to create a new error type for a given example, not yet present in a pool. Now you need to generate SHORT LABEL and a 1 or 2 SENTENCE DESCRIPTION for this new error type. Please try to be very specific in determining a FINE GRAINED error type, avoid too much generic error types. At the same time, it is important that the generated label and description of the error type can be GENERALIZED to other examples, i. e. avoid references to the particular content of the current example (names, dates, etc): anything related ONLY to this example SHOULD NOT be present in the description and label.
Output answer is the following format: "SHORT LABEL: DESCRIPTION", do not output anything else!
Figure 8: Prompt used for generating a new issue type in the cumulative clustering strategy. The presented version of the prompt is for text LLM outputs, the prompt can be easily changed if JSON outputs are supported by an LLM.
Situation: Two experts are inspecting examples in natural language generation for a particular dataset.
You will be given 2 sentences, which represent the conclusions of the two experts about the same example, i. e. what is a failure case in the example they were given.
Your task is to determine if the experts describe the same failure.
Expert 1 conclusion: E1
Expert 2 conclusion: E2
Do the experts describe the same failure? Output only one word ‘Yes’ or ‘No’.
Figure 9: Prompt used for evaluation. The presented version of the prompt is for text LLM outputs, the prompt can be easily changed if JSON outputs are supported by an LLM.
Appendix B Clustering experiment setup
In this experiment, we perform a hyperparameter grid search for five clustering algorithms: KMeans, Agglomerative Clustering, Spectral Clustering, Gaussian Mixture Models (GMM), and HDBSCAN on a synthetic set. Each algorithm is evaluated across a range of hyperparameter combinations. For KMeans, we vary the distance_metric (euclidean, cosine), kmeans_init strategy (k-means++, random), kmeans_n_init (10, 50), and kmeans_max_iter (300, 500). For Agglomerative Clustering, we test all combinations of distance_metric (euclidean, cosine) and linkage_type (ward, average, complete), while ensuring that ward is only paired with euclidean (as required by the algorithm). Spectral Clustering configurations include distance_metric (euclidean, cosine), assign_labels (kmeans, discretize), spectral_gamma (0.1, 0.5, 1.0, 2.0), and spectral_n_neighbors (5, 10, 20). For GMM, we explore covariance_type (full, diag), gmm_init_params (kmeans, random), and gmm_max_iter (100, 300). Lastly, HDBSCAN is tested with distance_metric (euclidean, cosine), min_cluster_size (3, 5, 10, 15, 20), hdbscan_min_samples (None, 1, 5), and hdbscan_cluster_selection_method (eom, leaf). Each valid configuration is evaluated over three independent trials with different random seeds to ensure robustness. After collecting results based on Adjusted Rand Index (ARI), the best-performing configuration for each algorithm on the synthetic validation set is selected. These best configurations are then applied to the test set of synthetic data and to the real dataset.
Appendix C Models
Model BibTeX License Model Repository
GPT-4o OpenAI et al. (2024) Proprietary https://platform.openai.com/docs/models/gpt-4o
Gemini-2.0-Flash Anil et al. (2023) Proprietary https://deepmind.google/technologies/gemini/flash/
Qwen-2.5 Qwen et al. (2025) Apache 2.0 https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e
DeepSeek-R1-Distill-Llama-8B DeepSeek-AI (2025) Llama 3.1 Community License https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Aya-Expanse-8B Dang et al. (2024) Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/CohereLabs/aya-expanse-8b
Llama-3.1-8B-Instruct Grattafiori et al. (2024) Llama 3.1 Community License https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Llama-3.2-1B-Instruct Grattafiori et al. (2024) Llama 3.2 Community License https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct
Ministral-8B-Instruct Mistral AI (2024) Mistral AI Research License https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
Solar-10.7B Kim et al. (2024a) Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/upstage/SOLAR-10.7B-Instruct-v1.0
Vicuna-1.5-13B Chiang et al. (2023) Llama 2 Community License Agreement https://huggingface.co/lmsys/vicuna-13b-v1.5
Command-R-35B Gomez and for AI (2024) Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/CohereLabs/c4ai-command-r-v01
Table 5: References to the used LLMs; all LLMs allow use for research.
Appendix D Datasets
Dataset name Dataset reference License
Natural Language Generation
FLASK data mix: Self-Instruct, WizardLM, Koala, CommonSense QA Wang et al. (2023); Xu et al. (2024); Geng et al. (2023); Talmor et al. (2021) Apache 2.0, MIT, Apache 2.0, Creative Commons Attribution 4.0
WMT’22 Kocmi et al. (2022) Apache 2.0
Elitr-Bench Thonet et al. (2025) Attribution 4.0 International
PIZZA Arkoudas et al. (2022) Attribution-NonCommercial 4.0 International
GSM8K Cobbe et al. (2021) MIT License
ParaDetox Dementieva et al. (2024) OpenRAIL++
Retrieval-augmented QA
MKQA (ru) Longpre et al. (2021) Creative Commons Attribution-ShareAlike 3.0 Unported License
BioASQ Krithara et al. (2023) Attribution 2.5 Generic
RobustQA Santhanam et al. (2022); Han et al. (2024, 2023) Apache-2.0
SearchQA Dunn et al. (2017) BSD 3-Clause
SyllabusQA Fernandez et al. (2024) Attribution-NonCommercialShareAlike
Table 6: References to the used datasets; all datasets allow use for research. We select instances from test splits.
Appendix E Examples
The following pages present examples of task instances, per-instance analysis, generated error reports, and clustering confusion matrices, for all 12 considered datasets. Clusters of size 1 are shown in confusion matrices but omitted in error reports, for space purposes.
A zip archive with experimental results is available at https://github.com/tunde-ajayi/llm-as-a-qualitative-judge/tree/main/data.
E.1 Semantic parsing (PIZZA dataset)
E.2 Long context QA (Elitr-Bench dataset)
E.3 Detoxification (MultiDetox dataset)
E.4 Translation en-ru (WMT’22 dataset)
E.5 Instruction following (FLASK dataset)
E.6 Grade school math (GSM8K dataset)
E.7 Factoid QA in Russian (MKQA dataset)
E.8 Biomedical QA (BioASQ dataset)
E.9 Lifestyle forum QA (RobustQA dataset)
E.10 Writing forum QA (RobustQA dataset)
E.11 Search engine queries (SearchQA dataset)
E.12 Educational QA (SyllabusQA dataset)